# Zero-shot sampling of adversarial entities in biomedical question answering

### v1

@@ Abstract
The increasing depth of parametric domain knowledge in large language models (LLMs) is fueling their rapid deployment in real-world applications. In high-stakes and knowledge-intensive tasks, understanding model vulnerabilities is essential for quantifying the trustworthiness of model predictions and regulating their use. The recent discovery of named entities as adversarial examples in natural language processing tasks raises questions about their potential guises in other settings. Here, we propose a powerscaled distance-weighted sampling scheme in embedding space to discover diverse adversarial entities as distractors. We demonstrate its advantage over random sampling in adversarial question answering on biomedical topics. Our approach enables the exploration of different regions on the attack surface, which reveals two regimes of adversarial entities that markedly differ in their characteristics. Moreover, we show that the attacks successfully manipulate token-wise Shapley value explanations, which become deceptive in the adversarial setting. Our investigations illustrate the brittleness of domain knowledge in LLMs and reveal a shortcoming of standard evaluations for high-capacity models [^1] The code and datasets developed for the work will be made available on GitHub..
@@ 1 Introduction
LLMs such as pre-trained and finetuned generalist language models (GLMs) and their domain-adapted versions (Zhao et al., 2023; Wang et al., 2023a) incorporating transformer architectures are emerging as the technological backbone of next-generation search engines and knowledge bases (Petroni et al., 2019; Razniewski et al., 2021; Sung et al., 2021). They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Nat, 2023; Spitale et al., 2023; Si et al., 2023). Adversarial examples are valuable probes for model vulnerability and robustness, and are essential for performance evaluation and regulatory audits (Shayegani et al., 2023). For language models, adversarial examples are generated through guided search (Alzantot et al., 2018), sampling and probabilistic methods (Ren et al., 2019; Yang et al., 2020; Yan et al., 2022), or collected dynamically by human operators (Jia and Liang, 2017; Wallace et al., 2019b). While the search methods iteratively find random string adversarial examples (Wallace et al., 2019a), the sampling approach obviates the need for predetermined search heuristics. It can operate with only blackbox access to models and in the zero-shot setting, which is fitting for model auditing frameworks (Metaxa et al., 2021).
Figure 1: Text substitution attack on QA with adversarial distractors. (a) Illustration of the attack scheme in embedding space. A is the key to the question, B-D the original distractors, ${}^{\prime}$ and ${}^{\prime\prime}$ the adversarial distractors semantically close to and far away from the key A, respectively. (b) An example of the attack showing the correct (left) and incorrect (right) model responses (checkmarked) before and after perturbation.
@@
The present work focuses on question answering (QA) in the biomedical domain, which has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023a; Kung et al., 2023; Singhal et al., 2023b). GLMs have general biomedical knowledge during model development. BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts from scholarly publications (Tinn et al., 2023). Although QA is frequently employed for model performance evaluation (Clark and Etzioni, 2016; Gardner et al., 2019; Robinson and Wingate, 2023), currently, the performance of LLMs in biomedical QA lacks systematic evaluation and their deployment in the real world is still in its infancy (Kung et al., 2023). Specifically, model performance against adversarial examples is seldomly assessed and their efficient sampling in the biomedical domain remains challenging, leaving a gap in our understanding of model robustness in entity-rich, high-stakes domain.
Knowledge and reasoning tests in the biomedical domain assess the understanding of diverse concepts, which are represented by domain-specific named entities (NEs). In this work, we develop an in-domain adversarial probe for model knowledge on biomedical NEs through text substitution of entity mentions in the distractors of multiple-choice QA. We introduce a powerscaled distance-weighted sampling (DWS) method in embedding space to generate adversarial examples of distractors, or adversarial distractors, in the zero-shot setting. Our work reports the first use of DWS for adversarial attacks in text [^2] Throughout the text, we use DWS to refer to the class of sampling method that employs distance-dependent probability weights. The powerscaled DWS is a type of DWS, so is the other version introduced in (Wu et al., 2017).. The attack scheme enables traversing between the two regimes of adversarial entities semantically close to or far away from an anchor point (illustrated as the key A in Fig. 1a). For demonstration, we focus on attacking two of the most common NEs in the biomedical domain (Wei et al., 2019): drug and disease names, shown in an example in Fig. 1b.. In addition, we introduce the FDA-drugs and CTD-diseases datasets as the source vocabularies of the perturbation sets, along with an entity type-annotated version of existing biomedical QA datasets MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022). Through analysis of attack statistics, we show that our approach provides more flexibility to uniform random sampling in entity-centric attacks (Yan et al., 2022), which maps the entity-level robustness landscapes of different LLMs. We further explore the relationship between Shapley value-based model explanation (Chen et al., 2023) and adversarial attacks and uncover the defining signatures of successful attacks in this context.
@@ 2.1 Adversarial distractor generation
We focus on constructing adversarial distractors in multiple-choice QA. Let each question $\mathcal{Q}$ be represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$, a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$. The distractors are the wrong choices given to confuse test-takers, while the key is the correct one (CH and Saha, 2020). QA is a natural language generation task. In an autoregressive LLM, the answer corresponds to the next tokens generated following $(\mathcal{C},\mathcal{S},\mathcal{D},k)$. The adversarial distractor is a threat model in which the perturbed distractor set $\mathcal{D}^{\prime}$ takes the form
@@
$\mathrm{argmax}$ $\Pr(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)$
$\neq\mathrm{argmax}\Pr(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (2)
The formulation in Eq. (2) is an untargeted attack, while the targeted version only requires restricting the answer to the perturbed question to a specific distractor. We extend Eq. (2) to a more general case in Appendix A with multiple entities and additional text in the distractors. Unlike text addition attacks that introduce distracting information (Jia and Liang, 2017; Wallace et al., 2019a), the edit constraint on $\mathcal{Q}$ that is usually enforced in text adversarial attacks (Roth et al., 2021) doesn’t preclude large semantic perturbations at the entity level in the distractors. In this work, we draw perturbations from a finite perturbation set, denoted as -$\mathcal{E}_{\textrm{-}k}\coloneqq\mathcal{E}\setminus\{k\}$, with =$\mathcal{E}=\{e_{j}\}_{j=1}^{m+1}$ being the entity dataset as illustrated in Fig. 1a. In practice, $\mathcal{E}$ may be compiled from potential model use cases or curated by regulatory bodies in a model audit.
@@ 2.2 Powerscaled distance-weighted sampling
We consider sampling with distance-dependent probability weights from the embedding space of NEs. The concept of DWS was previously introduced in contrastive learning in computer vision, where the Euclidean distance between samples and an anchor point was used and the probability weights depend on the embedding dimensionality (Wu et al., 2017). We propose a more flexible version independent of the embedding dimensionality using a powerscaled distance with a tunable hyperparameter. Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}$ to entity $e_{j}$ in -$\mathcal{E}_{\textrm{-}k}$ such that the distractor $d$ is substituted by
@@
Here, $h(k,e_{j})$ is a distance metric between $k$ and $e_{j}$, while the exponent $n\in\mathbb{R}$ controls the shape of the probability weight distribution. When $n>0$, as $n$ increases, the entities further away from $k$ are more preferably weighted. When $n<0$, the sampling method employs inverse distance weights, the entities closer to $k$ are more preferably weighted. Besides, the formalism accommodates a deterministic regime as a limiting case [^3] The limit $n\rightarrow-\infty$ indicates replacing the distractor $d$ with $k$ at all times. Since $k$ is not in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this limit is not attained.. It corresponds to $n\rightarrow+\infty$, when $p_{j}$ is vanishingly small for all but the entity furthest from $k$ in the embedding space according to $h(k,e_{j})$,
@@
Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq. (3)) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks. The observed similar behaviors are compared across models and datasets in (a) Flan-T5-XXL on MedQA-USMLE, (b) MedAlpaca-7B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA. Disease and drug-mention questions are separated by colors. The vertical dashed line indicates the location of random sampling.
3 Datasets

### v2

# Assessing biomedical knowledge robustness in large language models by query-efficient sampling attacks

@@ Abstract
The increasing depth of parametric domain knowledge in large language models (LLMs) is fueling their rapid deployment in real-world applications. Understanding model vulnerabilities in high-stakes and knowledge-intensive tasks is essential for quantifying the trustworthiness of model predictions and regulating their use. The recent discovery of named entities as adversarial examples (i.e. adversarial entities) in natural language processing tasks raises questions about their potential impact on the knowledge robustness of pre-trained and finetuned LLMs in high-stakes and specialized domains. We examined the use of type-consistent entity substitution as a template for collecting adversarial entities for billion-parameter LLMs with biomedical knowledge. To this end, we developed an embedding-space attack based on powerscaled distance-weighted sampling to assess the robustness of their biomedical knowledge with a low query budget and controllable coverage. Our method has favorable query efficiency and scaling over alternative approaches based on random sampling and blackbox gradient-guided search, which we demonstrated for adversarial distractor generation in biomedical question answering. Subsequent failure mode analysis uncovered two regimes of adversarial entities on the attack surface with distinct characteristics and we showed that entity substitution attacks can manipulate token-wise Shapley value explanations, which become deceptive in this setting. Our approach complements standard evaluations for high-capacity models and the results highlight the brittleness of domain knowledge in LLMs [^1] The code and datasets developed for the work is available at https://github.com/RealPolitiX/qstab/..
@@ 1 Introduction
LLMs such as pre-trained and finetuned generalist language models (GLMs) and their domain-adapted versions (Zhao et al., 2023; Wang et al., 2023a) incorporating transformer architectures are emerging as the technological backbone of next-generation search engines and knowledge bases (Petroni et al., 2019; Sung et al., 2021). They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Anonymous, 2023; Spitale et al., 2023; Si et al., 2023). Adversarial examples are valuable probes for model vulnerability and robustness, and are essential for performance improvements and regulatory audits (Shayegani et al., 2023). Adversarial attacks are classified by query scaling behavior into low- and high-query-budget attacks (see Fig. 1a). Collecting adversarial data at scale, either for adversarial training or robustness assessment, requires query-efficient attacks, first empirically studied in the vision domain (Ilyas et al., 2018; Cheng et al., 2019; Shukla et al., 2021). For language models, adversarial examples are generated through guided search (Alzantot et al., 2018), sampling and probabilistic methods (Ren et al., 2019; Yang et al., 2020; Yan et al., 2022), or collected dynamically by human operators (Jia & Liang, 2017; Wallace et al., 2019b). Most of these methods require a large amount of model evaluations (in the hundreds to thousands or beyond per instance) or whitebox access to model internals, and their evaluations tend to focus on small models (with < 1B parameters) (Maheshwary et al., 2021), which are insufficient in the era of LLMs. Sampling-based attacks obviate the need for predetermined search heuristics and can operate with only blackbox model access and a low query budget. Demonstrations of query-efficient attacks on LLMs meet an emerging demand for their continuous monitoring (Metaxa et al., 2021), which is particularly important for deploying LLMs in specialized domains.
@@
Question answering (QA) has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023a; Kung et al., 2023; Singhal et al., 2023b). Knowledge and reasoning components in biomedical QA assess the understanding of diverse concepts, which are represented by domain-specific named entities (NEs). The present work focuses on the identification of adversarial entities in biomedical QA for both GLMs and BLMs. While GLMs have general biomedical knowledge during model development, BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts from scholarly publications (Tinn et al., 2023). Although QA is frequently employed paradigm for model performance evaluation (Gardner et al., 2019; Robinson & Wingate, 2023), currently, the performance evaluation in biomedical QA lacks robustness assessment and their deployment in the real world, albeit promising, is still in its infancy (Kung et al., 2023). Analyzing biomedical texts using LLMs requires them to have robust contextual knowledge and language understanding of entities composed of single or multiple words. Evaluating model robustness in prediction for high-stakes, specialized domains requires customized methods apart from traditional benchmarks.
(a) (a)
(b) (c)
(c) (b)
Figure 1: Text substitution attack on QA with adversarial distractors. (a) Typical query scaling curves in low- and high-query-budget attack settings. (b) An adversarial distractor example found by type-consistent entity substitution (highlighted in blue). The correct (top) and incorrect (bottom) model responses (checkmarked) before and after perturbation are included. (c) Illustration of the attack scheme in embedding space by PDWS for the example in (b). $\mathcal{E}$ represents the vocabulary set. D is the key to the question, A-C the original distractors, C is an adversarial distractor at distance $h$ from the center at key D. The locations of A-C and C are relative to D in the embedding space.
In this work, we developed an effective in-domain, query-efficient adversarial probe for the knowledge robustness of LLMs in entity-rich domains using type-consistent entity substitution (TCES). They operate when a large number of viable replacements ($N_{\mathcal{E}}$ > 1000) per attack exist, while only a small number of text locations are permissible. The setting is prevalent in specialized domains where the NEs have shared characteristics such as their type and (word or phrase) structure. It differs from traditional synonym substitution attacks (Liu et al., 2022), which use a much smaller number of viable replacements per attack with matching word senses but can operate on many potential text locations. Our focus is on generating adversarial examples of distractors (see Fig. 1b), or adversarial distractors in biomedical QA tasks. For relevance to real-world scenarios (Apruzzese et al., 2023), we consider the attacks in the blackbox and hard-label setting with a limited query budget ($B\ll N_{\mathcal{E}}$). Our main contributions include: (i) Introduction of powerscaled distance-weighted sampling (PDWS) method in embedding space to generate adversarial entities. Our work reports the first use of DWS for adversarial attacks in text [^2] Throughout the text, we use DWS to refer to the class of sampling method that employs distance-dependent probability weights. The powerscaled DWS is a type of DWS, so is the other version introduced in (Wu et al., 2017).. The attack scheme has a broad coverage of the embedding space centered around an anchor point (illustrated as the key D in Fig. 1c) and maps the entity-level robustness landscapes of different LLMs in two regimes. (ii) Introduction of the FDA-drugs and CTD-diseases datasets as the source vocabularies of the perturbation sets, along with an entity type-annotated version of existing biomedical QA datasets MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022). Using these data, we demonstrated the vulnerability of LLMs to two of the most common NEs in the biomedical domain (Wei et al., 2019): drug and disease names (see example in Fig. 1b). (iii) We unified sampling- and gradient search-based methods in TCES as an attack template, which we used for quantitative comparison of the scaling behaviors of the attacks in the same footing. We found that in the blackbox setting, sampling attacks can have favorable query scaling over gradient-based attacks, when are limited by the quality of gradient estimation within a fixed query budget. (iv) By investigating the relationship between Shapley value-based model explanation (Chen et al., 2023) and adversarial attacks, we uncovered the defining signatures of successful attacks as a result of entity substitution.
2 Related works
Text substitution attack
Adversarial attacks in neural language models have been demonstrated from the character to the sentence level (Zhang et al., 2020). The majority of text substitution attacks employ typos (Gao et al., 2018) and synonyms (Mozes et al., 2021), which effectively impose lexical constraints. Word frequency has been used as a metric to select adversarial examples in (Mozes et al., 2021). Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022). Alzantot et al. (2018) used a genetic algorithm to search for adversarial word substitution. Model sensitivity to entity-level perturbations such as substitution and swapping create problematic situations for knowledge-intensive tasks, including knowledge conflicts (Longpre et al., 2021), compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
Robustness in QA
The robustness of QA models has been studied using perturbation methods (Jia & Liang, 2017) and meaning-preserving transformations (Gan & Ng, 2019; Elazar et al., 2021). Sen & Saffari (2020) introduced a set of perturbations to assess language model generalization between QA datasets. Richardson & Sabharwal (2020) proposed a distractor design method using proximity in semantic space to probe model understanding of word senses. Awadalla et al. (2022) examined the connection between in- and out-of-distribution robustness for different LLMs. The performance of LLMs is sensitive to the ordering of options (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024). Overall, the brittleness of knowledge in LLMs remains a primary limitation for their applications (Elazar et al., 2021; Augenstein et al., 2024).
Adversarial attack characteristics
The defining characteristic of adversarial attacks is to induce false predictions in trained models. Other characteristics include (i) their generality such as transferability, universality (Zou et al., 2023). Transferability refers to the effectiveness of the same example being adversarial to distinct models (Papernot et al., 2016). Universal adversarial perturbations are reusable across many examples tested on the same model (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a). (ii) The quality of adversarial examples may be judged by their perceptibility to human (Dyrmishi et al., 2023). (iii) Adversarial attacks can operate in the low- and high-query-budget settings (Shukla et al., 2021). (iv) Attacked text using specifically designed perturbations to the inputs can obstruct model explanation (Noppel & Wressnegger, 2024) and render it unuseful.
@@ 3.1 Adversarial distractor generation
We focus on constructing adversarial distractors in multiple-choice QA. Let each question $\mathcal{Q}$ be represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$ [^3] The stem in the example of Fig. 1b is the last sentence starting with “which” on the left side. The context is the text before the stem., a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$. The distractors are the wrong choices given to confuse test-takers, while the key is the correct one (CH & Saha, 2020). QA is a natural language generation task. In an autoregressive LLM, the answer corresponds to the next tokens generated following $(\mathcal{C},\mathcal{S},\mathcal{D},k)$. The adversarial distractor is a threat model in which the perturbed distractor set $\mathcal{D}^{\prime}$ takes the form
@@
$\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)\neq\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (2)
The formulation in Eq. 2 is an untargeted attack, while the targeted version only requires restricting the answer to the perturbed question to a specific distractor. We extend Eq. 2 to a more general case in Appendix A with multiple entities and additional text in the distractors. Unlike text addition attacks that introduce distracting information (Jia & Liang, 2017; Wallace et al., 2019a), the edit constraint on $\mathcal{Q}$ that is usually enforced in text adversarial attacks (Roth et al., 2024) doesn’t preclude large semantic perturbations at the entity level in the distractors. In this work, we draw perturbations from a finite perturbation set, denoted as -$\mathcal{E}_{\textrm{-}k}\coloneqq\mathcal{E}\setminus\{k\}$, with =$\mathcal{E}=\{e_{j}\}_{j=1}^{m+1}$ being the entity dataset as illustrated in Fig. 1a. In practice, $\mathcal{E}$ may be compiled from potential model use cases or curated by regulatory bodies in a model audit.
@@ 3.2 Powerscaled distance-weighted sampling
We consider sampling with distance-dependent probability weights from the embedding space of NEs. The concept of distance-weighted sampling (DWS) was previously introduced in contrastive learning in computer vision, where the Euclidean distance between samples and an anchor point was used and the probability weights depend on the embedding dimensionality (Wu et al., 2017). We propose a more flexible version independent of the embedding dimensionality using a powerscaled distance with a tunable hyperparameter. Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}$ to entity $e_{j}$ such that the distractor $d$ is substituted by
@@
Here, $h(k,e_{j})$ is a distance metric between the anchor point $k$ and $e_{j}$, while the exponent $n\in\mathbb{R}$ controls the shape of the probability weight distribution. When $n>0$, as $n$ increases, the entities further away from $k$ are more preferably weighted. When $n<0$, the sampling method employs inverse distance weights, and the entities closer to $k$ are more preferably weighted. In the subsequent evaluations, we use $h(k,e_{j})=\texttt{CosineDist}(\mathbf{k},\mathbf{e}_{j})=1-\mathbf{k}\cdot\mathbf{e}_{j}/(\|\mathbf{k}\|\cdot\|\mathbf{e}_{j}\|)$, calculated in a text embedding of choice. We use here the italic letters ($k,e$) to represent text while the bolded letters ($\mathbf{k},\mathbf{e}$) are their corresponding vector representations. The formalism accommodates a deterministic regime as a limiting case [^4] The limit $n\rightarrow-\infty$ indicates replacing the distractor $d$ with $k$ at all times. Since $k$ is not in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this limit is not attained.. It corresponds to $n\rightarrow+\infty$, when $p_{j}$ is vanishingly small for all but the entity furthest from $k$,
@@
3.3 Sampling in zeroth-order gradient attack
Adversarial attacks based on zeroth-order optimization (ZOO) use estimated gradients from finite difference to guide the search (Ilyas et al., 2018; Cheng et al., 2019). In the discrete text space, Berger et al. (2021) developed the DiscreteZOO attack, which contains three components: the word importance ranker, the candidate sampler, and the gradient-based optimizer. It operates on the word level and is validated on BERT-sized models for synonym substitution attacks. We discuss here the last two components, which, in each iteration, aim to find a replacement text span $e^{\prime}$ through the gradient update rule
$\mathbf{e}^{\prime}=\mathbf{e}_{0}-\lambda\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0}),$ (5)
where $e_{0}$ is the original text span, $\lambda$ the learning rate, and ^$\widehat{\nabla}_{\mathrm{ZO}}$ indicates the zeroth-order gradient estimator. We use $\mathrm{LLM}(\mathcal{Q};e_{0})$ to denote the LLM output with the unperturbed question $\mathcal{Q}$ (containing $e_{0}$) as the input, then the multi-point ($M\geq 2$) gradient estimate is constructed by querying the LLM $M$ times and computing
^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})=\mathbb{E}_{\textbf{u}}\left[\frac{\Delta\mathrm{LLM}(\mathcal{Q};e_{0})}{\Delta\mathbf{e}_{0}}\textbf{u}\right]=\frac{1}{M}\sum_{i=1,e_{i}\in\mathcal{E}_{\textrm{-}k}}^{M}\frac{\left(\mathrm{LLM}(\mathcal{Q^{\prime}};e_{i})-\mathrm{LLM}(\mathcal{Q};e_{0})\right)\cdot(\mathbf{e}_{i}-\mathbf{e}_{0})}{\|\mathbf{e}_{i}-\mathbf{e}_{0}\|^{2}}.$ (6)
This is an example of the random directions stochastic approximation (Nesterov & Spokoiny, 2017), where $\textbf{u}=(\mathbf{e}_{j}-\mathbf{e}_{0})/\|\mathbf{e}_{j}-\mathbf{e}_{0}\|$ is a randomly oriented vector in the embedding space corresponding to the neighboring entity $e_{j}$. Berger et al. (2021) computed Eq. 6 by random sampling of neighbors (in candidate sampler) within a similarity threshold for synonyms. Given a noisy gradient estimate ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})$ and the discrete nature of the search space, we can rewrite Eq. 5 effectively as deterministic DWS in the embedding space by
$d^{\prime}=e_{j}\quad\mathrm{s.t.}\quad e_{j}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmin}}\,h(e_{0},e_{j}),$ (7)
where $h(e_{0},e_{j})=\texttt{CosineDist}(\mathbf{e}_{0}-\lambda\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0}),\mathbf{e}_{j})$, with $e_{0}$ as the anchor point, and argmin indicates the closest member within the text embedding. Using argmin is essential because the exact location $h(e_{0},e_{j})$ in the discrete embedding space is usually not occupied. While a gradient estimate of sufficient quality is essential for a successful attack, increasing $M$ will count towards the query budget. Therefore, in the budgeted setting, a tradeoff should exist between the efficacy of the attack and the number of queries consumed for each gradient estimation.
4 Implementing attacks on biomedical question answering
4.1 Dataset selection